OpenStudy (anonymous):

What is 7 · 2 and 3/7?

OpenStudy (calculusxy):

What are you trying to say by that?

OpenStudy (anonymous):

What is the answer to 7 multiplied by 2 and 3/7ths?

OpenStudy (jdoe0001):

\(\bf 7\times 2\cfrac{3}{7}\quad ?\)

OpenStudy (anonymous):

No, in fraction form

OpenStudy (anonymous):

and @jdoe0001 yes

OpenStudy (jdoe0001):

\(\bf 7\times 2\cfrac{3}{7}\implies 7\times \cfrac{2\times 7 +3}{7}\implies 7\times\cfrac{17}{7}\implies \cfrac{\cancel{7}\times 17}{\cancel{7}}\)